OpenStudy (anonymous):

ans of hw 10 pls??????

OpenStudy (anonymous):

H10P1 Circuit A: r*u Circuit B: q*v Circuit C: p*w Circuit D: s*t H10P2 A) R1+R2+j*w*L+(1/(j*w*C)) B) inf C) inf D) ((((R1+R2)*(j*w*L))/((R1+R2)+(j*w*L)))*(1/(j*w*C)))/((((R1+R2)*(j*w*L))/((R1+R2)+(j*w*L)))+(1/(j*w*C))) E) 0 F) 0 G) ((R1+(1/(j*w*C)))*(R2+j*w*L))/((R1+(1/(j*w*C)))+(R2+j*w*L)) E) R2 F) R1 H10P3 a) (((j*w*L+R2)*(1/(j*w*C))/(j*w*L+R2+(1/(j*w*C))))/((j*w*L + R2)*(1/(j*w*C))/(j*w*L+R2+(1/(j*w*C)))+R1)*(R2/(j*w*L + R2))) b) ((j * w * L + R2) * (1 / (j * w * C))) / (j * w * L + R2 + (1 / (j * w * C ))) c) (sqrt (R1/R2 -1)) / (w * R1) d) ((sqrt (R1/R2 -1)) * R2) / w e) Change values in C (with w = 2*pi*f) be careful, because the f is normaly in Khz (*10^3) f) Change values in D (with w = 2*pi*f) be careful, because the f is normaly in Khz (*10^3)

OpenStudy (anonymous):

Last two I can't get them right :( so, following the C formula for e): for R1= 1000 , R2 = 50 and f=870kHz ((sqrt (1000/50 -1))/(2*pi*870k*1000) = 7.97402658968e-10 it wont check still :/

OpenStudy (anonymous):

capacitance (in picoFarads) and inductance (in microHenrys) 7.97402658968e-10 = 797.402658968pF Try 797.402658968

OpenStudy (anonymous):

Thank you so much! it worked! :D

OpenStudy (anonymous):

You are welcome